OpenStudy (anonymous):

A wall is made of both red and yellow bricks. The ratio of red bricks to yellow bricks 5:2. If there are 210 more red brick then yellow, how many bricks are in the wall?

OpenStudy (anonymous):

For every 5 red bricks there are 2 yellow bricks. Yellow = X Red = X + 210 Since the red:yellow ratio doesn't change, (5/2) = (210+X)/X Cross multiplying gets you 5X = 2(210+X). Solve for X. Once you know how many yellow and red bricks there are, add them up and you have the total.

OpenStudy (anonymous):

why did you put the red bricks/yellow bricks?

OpenStudy (anonymous):

It could be the other way around, it's just a ratio. You could also say instead of a 5:2 ratio it's 2.5:1 ratio. Then for X yellow bricks you have 2.5X red bricks. 2.5X = 210 + X. If you solve for the original equation: 5X = 420 + 2X 2(2.5X) = 2(210 + X) Same equation.